OpenStudy (anonymous):

given zero to find the remaining zeros of the function f(x)=x^3-8x^2+25x-200; zero -5i

OpenStudy (anonymous):

Hint:\[x^3-8 x^2+25 x-200=(x-8) \left(x^2+25\right) \]